# Disentangle and Regularize: Sign Language Production with Articulator-Based Disentanglement and Channel-Aware Regularization

### v2

@@ Abstract
In this work, we propose DARSLP, a simple gloss-free, transformer-based sign language production (SLP) framework that directly maps spoken-language text to sign pose sequences. We first train a pose autoencoder that encodes sign poses into a compact latent space using an articulator-based disentanglement strategy, where features corresponding to the face, right hand, left hand, and body are modeled separately to promote structured and interpretable representation learning. Next, a non-autoregressive transformer decoder is trained to predict these latent representations from sentence-level text embeddings. To guide this process, we apply channel-aware regularization by aligning predicted latent distributions with priors extracted from the ground-truth encodings using a KL-divergence loss. The contribution of each channel to the loss is weighted according to its associated articulator region, enabling the model to account for the relative importance of different articulators during training. Our approach does not rely on gloss supervision or pretrained models, and achieves state-of-the-art results on the PHOENIX14T and CSL-Daily datasets.
@@ 1 Introduction
Sign language serves as the primary mode of communication for Deaf and Hard-of-Hearing (DHH) individuals, relying on a rich combination of manual (hand movements) and non-manual (facial expressions, body posture) features. With the rise of deep learning, SLP has gained attention as a task that aims to generate sign language sequences from spoken language text. However, producing realistic and continuous sign language sequences remains a challenging problem due to the complex spatial-temporal nature of sign articulation.
Recent advances in SLP have focused on generating skeletal pose sequences instead of synthesizing sign language videos directly. This intermediate representation provides a structured and computationally efficient way to model sign articulation while separating the challenges of motion synthesis and realistic rendering [19]. Despite progress, generating natural and expressive sign pose sequences remains an open challenge. Many SLP models tend to produce averaged movements, failing to capture the fine-grained articulations required for natural sign language expressions [30], [17]. This results in overly simplified gestures, particularly in hand and finger movements. Additionally, error accumulation in autoregressive decoding, where errors propagate over time, further degrades sequence coherence, leading to unnatural transitions between signs [20] , [21]. To address these challenges, various techniques have been explored, including adversarial training, progressive transformers, and mixture density networks (MDNs) [22]. While these methods improve realism, they struggle with smooth and expressive sign transitions, particularly in fine hand articulations.
A significant research direction in SLP involves reducing dependency on gloss annotations, which traditionally serve as an intermediate representation between spoken and sign language. While gloss-based models achieve high accuracy, gloss annotation is costly, time-consuming, and often inconsistent across datasets, making large-scale deployment impractical. To overcome this limitation, gloss-free SLP has gained attention, aiming for direct text-to-sign generation without requiring gloss annotations [11, 30, 34]. While promising, these models still struggle with smooth transitions and expressive signing.
To address these challenges, we propose DARSLP, a simple transformer-based SLP framework that directly translates spoken language text into continuous sign pose sequences without gloss supervision. The contribution of our work to the SLP domain can be summarized as follows:
@@
  • We propose a channel-aware regularization strategy for training the non-autoregressive transformer by aligning predicted latent distributions with articulator-weighted channel priors derived from the autoencoder. In contrast to common variational methods that impose prior constraints, our approach directly computes these priors from observed channel distributions. Empirical results show that this regularization improves motion diversity and realism, and effectively reduces the regression-to-the-mean problem in generated pose sequences.
Unlike previous approaches that rely on gloss supervision, our model removes the need for gloss annotations while surpassing both gloss-based and gloss-free models in back-translation performance, achieving state-of-the-art results. Our model ranked third in the CVPR 2025 SLRTP Challenge [31], demonstrating its effectiveness in producing accurate pose generation in sign language.
@@ 2 Related Work
Sign language production aims to generate sign language sequences corresponding to a given spoken language text. Recently, studies have focused on converting input text into skeletal pose sequences instead of directly synthesizing sign language videos, as this intermediate representation is more practical and generalizable [19]. The synthesis of realistic videos from these pose sequences is considered a separate research challenge. In our study, we also concentrate on generating realistic pose sequences.
One of the primary challenges in continuous SLP, akin to continuous sign language translation (SLT), stems from the lack of one-to-one alignment between spoken language and sign language sequences. Producing natural and continuous sign language sequences is significantly more complex than simply concatenating isolated signs. Ensuring a smooth transition between signs while preserving semantic integrity is crucial for achieving realistic sign flow. However, existing SLP models struggle to model these transitions effectively.
@@ 2.1 Autoregressive Approaches
A major unsolved problem in SLP research is regression to the mean, a phenomenon in which generated sign sequences converge to average motion patterns, resulting in unnatural and overly simplified gestures [30], [17]. This issue reduces the expressiveness of generated signs, particularly in hand and finger movements, leading to deviations from the intended semantic content. Consequently, sign language users may find the generated sequences difficult to understand, lowering communication quality. Addressing this fundamental issue is essential to improve the realism and effectiveness of SLP systems.
The field of SLP has seen significant advancements since the pioneering work of Stoll et al. (2018), who used a neural machine translation (NMT) model to convert text to gloss sequences [23]. However, these early approaches were insufficient for producing natural and continuous sign language sequences. Later, Saunders et al. (2020a) introduced a transformer-based continuous SLP model. Despite improvements, the autoregressive nature of transformer-based models exacerbated the regression-to-the-mean issue, limiting their ability to capture fine details [20]. To mitigate this, Saunders et al. (2020b) incorporated adversarial training with a multichannel sign production framework. Although this approach systematically addressed the regression-to-the-mean problem for the first time, it did not fully resolve it [21].
Later methods such as progressive transformers and mixture density networks (MDN) have been explored, combined with data augmentation and adversarial training techniques [22]. Despite these efforts, user evaluations indicate that the regression-to-the-mean problem persists. While MDN models attempt to model the natural variation in sign sequences, they fail to capture fine-grained finger movements, resulting in uncorrect articulation patterns.
@@ 2.2 Non-Autoregressive Approaches
To overcome these limitations autoregressive approaches, Hwang et al. (2021) introduced the NSLP-G model [9]. This model encoded signs in a Gaussian space using a VAE framework to generate poses directly. However, NSLP-G struggled to predict sequence length and failed to capture fine-grained finger details effectively. Hwang et al. (2022) later proposed length regulators and knowledge distillation techniques to address these issues [10]. By replacing MSE loss with binary cross-entropy (BCE), they improved the capture of low-variance details. However, the challenge of representing fine hand and finger articulations in the latent space remains unresolved. The VAE encoder, despite producing distinct latent vectors, lacks the capacity to fully encode the intricacies of hand poses. The decoder tends to focus on reconstructing larger body structures while neglecting hand details.
Ma et al. (2024) tackled the regression-to-the-mean issue by introducing a dual-decoder transformer, designed to improve hand modeling in SLP. Their approach employed a separate decoder for manual features (hand poses) alongside the full-body pose decoder to ensure better alignment. While this method showed improvements in hand articulation, it also revealed critical challenges. The model’s performance suffered due to the scarcity of high-quality training data, particularly for complex hand poses. Moreover, the dual-decoder structure increased computational costs and was highly sensitive to dataset quality, limiting its generalizability [17].
@@ 2.3 Gloss-Free Approaches
While gloss-based SLP models achieve higher accuracy, gloss annotation is costly and time-consuming, limiting scalability. To address this, gloss-free approaches aim to translate spoken language directly into sign language sequences without gloss labels. Hwang et al. (2024) introduced SignVQNet [11], which employs an autoregressive structure by converting sign pose sequences into discrete tokens. Similarly, Walsh et al. (2024) leveraged vector quantization (VQ) to reduce gloss dependency. However, VQ-based discrete tokenization often results in unnatural transitions and loss of fine details in complex sign motions [30, 29]. Yin et al. (2024) proposed Dynamic Vector Quantization (DVQ-VAE), which adjusts encoding length based on information density [34]. Despite improvements, VQ-based methods still struggle with smooth sign transitions, and error accumulation in autoregressive models remains a challenge.
To alleviate the limitations posed by gloss annotations, several alternative intermediate representations have been proposed. Systems such as the Hamburg Notation System (HamNoSys) [7] and SignWriting [24] aim to capture phonetic or visual aspects of sign language and have been explored as substitutes for glosses in sign language production pipelines. For instance, HamNoSys, which encodes signs at the articulatory level, can be directly mapped to avatars and has been utilized in the Text-to-HamNoSys (T2H) task, showing improved translation performance [28]. Similarly, recent work has investigated SignWriting-based translation and synthesis pipelines [1, 14]. However, these systems still rely on specialized, manually annotated symbolic representations and suffer from similar scalability challenges as gloss-based models.
Given the data scarcity and annotation bottlenecks in sign language datasets our work follows an alternative paradigm by completely bypassing gloss-level supervision. We propose a gloss-free sign language production framework that learns to generate continuous sign pose sequences directly from spoken language inputs, enabling scalable and end-to-end learning without the need for intermediate linguistic annotations.
@@ 2.4 Latent Space Modeling and Disentanglement
Beyond discrete tokenization, some studies explored continuous latent space modeling, often using autoencoders or variational methods [9, 10]. These works attempt to learn compact, expressive embeddings that preserve both manual and non-manual features. However, they often treat pose data holistically, failing to separate the dynamics of distinct articulators. This design leads to inefficient representation of subtle articulations, particularly for hands and fingers. Since these models operate in a Gaussian latent space, they tend to capture global patterns while underrepresenting low-variance, fine-grained features.
Beyond the SLP domain, significant advancements have been made in learning compact, structured latent spaces using autoencoders across diverse modalities. Autoencoders have proven effective in capturing essential data representations while reducing dimensionality, supporting tasks such as sequence modeling [4], biomedical data analysis [32], and latent variable modeling [15]. Particularly in multimodal tasks, disentangled representations are crucial to separately model different modalities’ distinct characteristics while maintaining coherence. Inspired by this, our approach incorporates a structurally disentangled autoencoder specifically tailored to the multimodal nature of sign language, where manual (hand movements) and non-manual (facial expressions, body posture) features must be modeled jointly but distinctly. Unlike prior SLP works that treat pose sequences holistically or rely heavily on discrete representations, we structure the latent space to explicitly capture the separate dynamics of different body parts.
Additionally, while KL divergence loss has been widely employed in variational autoencoders (VAEs) and sequence generation tasks to regularize latent spaces [2], its potential remains underutilized in sign language production. In this work, we investigate the usage of KL divergence as a regularization signal within a structurally disentangled latent space, designed to represent manual and non-manual articulators separately. This setup allows for targeted regularization that promotes smoothness and structural consistency across body regions, without relying on a variational framework.
3 Metodology
@@ 3.1 Pose Autoencoder
Figure 1: Overview of the proposed pose autoencoder architecture.
@@
To effectively capture the multimodal structure inherent in sign language articulation, we employ a structurally disentangled pose autoencoder designed to encode and reconstruct sign poses in a compact latent space. Rather than aiming to achieve disentanglement in the strict unsupervised sense, we explicitly incorporated an architectural design that encourages the model to learn semantically organized representations. This design introduces a multi-modal latent factorization that aligns well with the physical structure and functional roles of each body part in sign language expression. The model operates on 3D skeleton sequences from the normalized version of the PHOENIX2014T dataset. Each input sequence consists of 50 upper body and hand joints alongside 128 facial keypoints, where each keypoint is represented by its (x, y, z) coordinates.
The encoder architecture shown in Figure 1 is explicitly partitioned to process four anatomically distinct regions: upper body, right hand, left hand, and face. Each region is assigned to a dedicated linear encoder. The latent representation is an 80-dimensional vector, within which the face component is fixed at 16 dimensions, and the remaining 64 dimensions are distributed between the upper body and hands. The latent dimensionality is allocated proportionally based on the number of joints in each region, ensuring balanced representation. The upper body region, consisting of 8 joints, corresponding to shoulders, elbows, wrists and hips, is mapped from a 8×3 vector to an 8-dimensional latent representation. The right and left hands, each containing 21 joints, are projected from 21×3 vectors to 28-dimensional latent vectors. The face region, comprising 128×3 facial landmarks, is compressed to a 16-dimensional latent space. These configurations are selected based on ablation studies during autoencoder training, where various latent sizes are compared in terms of pose reconstruction quality. Although a higher-dimensional space yielded slightly better performance, we selected the 80-dimensional configuration to maintain a compact latent representation, considering the limited size of training datasets typically available in the sign language domain.
The design choice to represent the face with only 16 dimensions, despite its high raw dimensionality, is grounded in our analysis of the keypoint variance for the landmarks, a significant portion of the facial keypoints exhibit highly correlated movement patterns, particularly in areas such as the cheeks, jawline, and forehead. These regions tend to move coherently during sign articulation, resulting in lower intrinsic variability. Consequently, the face keypoints form a manifold with lower effective dimensionality. By exploiting this redundancy, the encoder achieves a high compression rate without compromising reconstruction quality.
The decoder mirrors the encoder structure, reconstructing the original pose sequences from the structured latent vectors of each body region. To ensure high-fidelity reconstruction, we apply a weighted L1 loss, computing reconstruction losses separately for each region. To further enforce sparsity and improve the quality of the learned latent representations, we apply L1 regularization exclusively to the encoder weights which is computed as the sum of the absolute values of all encoder weight parameters. This regularization scheme ensures that the encoder learns compact and meaningful latent codes, which contributes to improved generalization and robustness in downstream sign generation.
@@ 3.2 Transformer Model
Following the latent pose representation learned by the pose autoencoder, the second stage of our framework generates sign pose embeddings conditioned on spoken language text. To accomplish this, we utilize a Transformer-based encoder-decoder architecture that directly maps sentence-level text embeddings into the structurally disentangled latent space previously established during the first stage of pose encoder training.
@@
The input to the transformer encoder consists of BERT-based text embeddings, where each word is represented by a 768-dimensional vector. These embeddings are projected to a 512-dimensional space through a linear transformation to align with the transformer’s internal dimensionality. The encoder comprises 3 layers, each equipped with 4 attention heads and a 1024-dimensional feed-forward network, followed by positional encoding to retain temporal ordering of the text sequence.
The decoder adopts a non-autoregressive structure to alleviate error accumulation issues commonly observed in autoregressive models. It includes 6 layers with 8 attention heads, each utilizing a 1024-dimensional feed-forward network. To initialize temporal dynamics, we employ learned time queries derived from a canonical reference pose, providing a consistent starting point for generating sign sequences. The decoder outputs are processed by a linear projection layer to reduce their dimensionality to match the 80-dimensional latent space defined by the autoencoder.
@@
Ground truth (GT) pose encodings for transformer training are obtained by passing the training set poses through the encoder of the pretrained pose autoencoder. These encodings provide compact, structurally disentangled representations of manual (hands) and non-manual (face, upper body) articulators, guiding the transformer model to learn structured mappings from spoken language text to continuous sign pose sequences. To enforce accurate reconstruction of these latent encodings, we apply L1 loss between the transformer’s predicted embeddings and the GT encodings.
Furthermore, to ensure that generated latent representations align with the distributional characteristics of the autoencoder’s latent space, we introduce a KL divergence-based regularization later in training. Specifically, prior distributions for each articulatory region, precomputed from the autoencoder, serve as reference distributions guiding the Transformer’s predictions. This approach promotes consistency and structural coherence in the generated embeddings, enhancing generalization to unseen inputs by maintaining fidelity to the learned latent space distribution.
@@
To align the duration of generated sign sequences with the complexity of the input sentences, we integrate a length predictor module into our framework. Specifically, the Transformer encoder outputs are pooled into a sentence-level embedding, which is then passed through a linear layer with sigmoid activation to predict a normalized length ratio between 0 and 1. This ratio corresponds to the proportion of the maximum possible sequence length required to represent the given input sentence. During inference, the predicted length ratio dynamically determines the actual sequence length by selecting the corresponding number of frames from the decoder’s output. The length predictor is trained using supervision from ground-truth sequence lengths, thereby promoting temporal coherence and improving the synchronization of generated sign motion with the input sentences. This also ensures that the rhythm and extent of sign motion appropriately reflect the sentence’s semantic load. This idea is further supported by previous work [34], which demonstrate a strong correlation between input sentence length and video duration, highlighting the importance of conditioning generation length on input characteristics.
@@
As shown in Figure 2, after the transformer generates the predicted pose embeddings, the sequence of 3D pose keypoints can be readily reconstructed by passing the output embeddings through the decoder of the pretrained pose autoencoder obtained in the first stage. This design enables efficient end-to-end generation of complete sign pose sequences without requiring additional refinement or post-processing steps.

### v3

@@ Abstract
In this work, we propose DARSLP, a simple gloss-free, transformer-based sign language production (SLP) framework that directly maps spoken-language text to sign pose sequences. We first train a pose autoencoder that encodes sign poses into a compact latent space using an articulator-based disentanglement strategy, where features corresponding to the face, right hand, left hand, and body are modeled separately to promote structured and interpretable representation learning. Next, a non-autoregressive transformer decoder is trained to predict these latent representations from word-level text embeddings of the input sentence. To guide this process, we apply channel-aware regularization by aligning predicted latent distributions with priors extracted from the ground-truth encodings using a KL divergence loss. The contribution of each channel to the loss is weighted according to its associated articulator region, enabling the model to account for the relative importance of different articulators during training. Our approach does not rely on gloss supervision or pretrained models, and achieves state-of-the-art results on the PHOENIX14T and CSL-Daily datasets.
@@ 1 Introduction
Sign language is the primary mode of communication for Deaf and Hard-of-Hearing (DHH) individuals, involving a rich interplay of manual (hand movements) and non-manual (facial expressions, body posture) features. With the rise of deep learning, sign language production (SLP)—the task of generating sign sequences from spoken text—has gained increasing attention. However, producing natural and continuous sign articulation remains challenging due to its complex spatial-temporal dynamics.
Recent efforts focus on generating skeletal pose sequences as an intermediate representation, which offers a structured and efficient way to model articulation, separating motion synthesis from video rendering [19]. Yet, generating natural and expressive sign sequences remains an open challenge. Many models regress to mean poses, losing detail in hand and finger movements [30, 17], resulting in overly simplified gestures. Autoregressive models often suffer from error accumulation, leading to unnatural transitions [20, 21]. To address these issues, techniques such as adversarial training, progressive transformers, and mixture density networks (MDNs) have been explored [22], but still fall short in producing smooth and expressive hand articulations.
A significant research direction in SLP is reducing reliance on gloss annotations, intermediate labels between spoken and sign language. Although gloss-based models achieve strong performance, gloss annotation is costly, time-consuming, and often inconsistent across datasets, limiting scalability. To address this, gloss-free SLP has gained attention, targeting direct text-to-sign generation without gloss supervision [11, 30, 35]. While promising, these models still face challenges in generating smooth transitions and expressive signing.
We propose DARSLP, a simple yet effective transformer-based SLP framework that directly translates spoken language text into continuous sign pose sequences without intermediate gloss annotations. The contribution of our work to the SLP domain can be summarized as follows:
@@
  • We propose a channel-aware regularization strategy for training the non-autoregressive transformer by aligning predicted latent distributions with articulator-based channel priors derived from the autoencoder. In contrast to common variational methods that impose prior constraints, our approach directly computes these priors from observed channel distributions. Empirical results show that this regularization improves motion diversity and realism, and effectively reduces the regression-to-the-mean problem in generated pose sequences.
Unlike most prior methods that depend on gloss supervision, our model eliminates the need for gloss annotations while surpassing both gloss-based and gloss-free baselines in back-translation performance. It achieved third place in the CVPR 2025 SLRTP Challenge [31], demonstrating state-of-the-art performance in sign pose generation.
@@ 2 Related Work
Sign language production (SLP) aims to generate sign sequences from spoken language text. Recently, studies have focused on producing skeletal pose sequences instead of full video synthesis, as this intermediate representation is more practical and generalizable [19]. Generating photorealistic videos from these poses remains a separate challenge. One of the primary challenges in continuous SLP, akin to continuous sign language translation (SLT), stems from the lack of one-to-one alignment between spoken language and sign language sequences. Realistic generation requires more than concatenating isolated signs; it demands smooth transitions and semantic coherence throughout the sequence. However, many SLP models still struggle to produce natural inter-sign dynamics.
@@ 2.1 Autoregressive Approaches
A major unsolved problem in SLP research is regression to the mean, a phenomenon in which generated sign sequences converge to average motion patterns, resulting in unnatural and overly simplified gestures [30], [17]. This leads to reduced expressiveness and semantic drift, lowering communication quality.
The field of SLP has seen significant advancements since the pioneering work of Stoll et al. (2018), who used neural machine translation (NMT) to map text to glosses, but failed to produce continuous, natural sign sequences [23]. Later, Saunders et al. (2020a) introduced a transformer-based SLP model, but its autoregressive decoding amplified the regression-to-the-mean effect [20]. To mitigate this, Saunders et al. (2020b) incorporated adversarial training with a multi-channel sign production framework. Although this approach systematically addressed the regression-to-the-mean problem for the first time, it did not fully resolve it [21].
Later, methods such as progressive transformers and mixture density networks (MDN) have been explored, combined with data augmentation and adversarial training techniques [22]. Despite these efforts, user evaluations indicate that the regression-to-the-mean problem persists. While MDN attempt to model the natural variation in sign sequences, it fails to capture fine-grained finger movements.
Our framework departs from autoregressive decoding entirely, eliminating sequential dependencies that amplify over-smoothing. By combining a non-autoregressive structure, which substantially reduces inference time [8, 10, 9], with a disentangled latent space, our approach addresses the regression-to-the-mean problem while producing expressive and natural motion without the need for adversarial objectives or complex multi-stage designs.
Figure 1: Overview of the proposed pose autoencoder architecture. Input 3D sign language poses are split into four regions: body (8 joints), right hand (21), left hand (21), and face (128). Each region is encoded separately to produce a compact 80-dimensional latent representation, then decoded by region-specific decoders to reconstruct the full pose sequence. A weighted L1 loss emphasizes critical articulators, such as the hands, during training.
@@ 2.2 Non-Autoregressive Approaches
To address limitations in autoregressive decoding, Hwang et al. (2021) proposed NSLP-G, which used a Variational Autoencoder (VAE) to generate sign poses in a Gaussian latent space [9]. However, it struggled with sequence length prediction and failed to capture fine-grained finger movements. A follow-up model introduced length regulators and replaced MSE with BCE loss, improving low-variance detail modeling [10]. Still, hand articulation remained poorly encoded, as VAEs tended to emphasize larger body motions and overlook fine manual features.
Ma et al. (2024) tackled the regression-to-the-mean problem by introducing a dual-decoder transformer, separating manual and full-body decoding [17]. Although this improved hand modeling, the approach suffered from high computational cost and sensitivity to data quality, limiting its scalability and generalization.
Our work advances non-autoregressive SLP by introducing a semantically structured latent space that explicitly disentangles hands, face, and body. This factorization allows for finer modeling of manual and non-manual features while maintaining computational efficiency. Furthermore, our length prediction mechanism provides robust temporal control without instability.
@@ 2.3 Gloss-Free Approaches
While gloss-based SLP models yield high accuracy, gloss annotation is costly and time-consuming, limiting scalability. To address this, gloss-free approaches aim to translate spoken language directly into sign language sequences without gloss labels. Hwang et al. (2024) proposed SignVQNet, converting pose sequences into discrete tokens within an autoregressive structure [11]. Walsh et al. (2024) also leveraged vector quantization (VQ) to reduce gloss dependency [30, 29]. However, VQ-based models often produce unnatural transitions and lose fine-grained motion detail. Yin et al. (2024) addressed this with Dynamic VQ (DVQ-VAE), which adapts encoding length to information density [35], but challenges with smooth transitions and autoregressive error accumulation persist.
Alternative symbolic systems like HamNoSys [7] and SignWriting [24] aim to replace glosses by encoding articulatory or visual features. These have shown promise in tasks like Text-to-HamNoSys (T2H) [28] and SignWriting-based synthesis [1, 14], but still require manual annotation and lack scalability.
In contrast, our work adopts a fully gloss-free approach that learns to generate continuous sign poses directly from spoken text, enabling end-to-end, scalable SLP without intermediate linguistic supervision.
@@ 2.4 Latent Space Modeling and Disentanglement
Beyond discrete tokenization, several studies have explored continuous latent space modeling using autoencoders or VAEs [9, 10]. These methods aim to learn compact embeddings capturing both manual and non-manual features. However, they typically treat pose data holistically, leading to inefficient encoding of fine-grained articulations, particularly for hands and fingers, while favoring global motion patterns.
Outside the SLP domain, autoencoders have demonstrated success in learning structured representations across modalities, supporting tasks such as sequence modeling [4], biomedical analysis [32], and latent variable modeling [15]. In multimodal settings, disentangling representations is key to modeling distinct signal sources while preserving coherence.
Motivated by this, we design a structurally disentangled autoencoder tailored to the multimodal nature of sign language, separating latent representations for hands, face, and body. Unlike prior SLP works that rely on holistic or discrete representations, our method explicitly captures the dynamics of each articulator group.
Additionally, we leverage KL divergence as a targeted regularization signal over these disentangled channels. While widely used in VAEs for latent control [2], KL loss remains underexplored in SLP. We apply it outside a variational setting to enforce structural coherence across regions, improving representation smoothness without sacrificing detail.
3 Methodology
Figure 2: Overview of the proposed DARSLP framework. Input text is encoded with a pre-trained BERT model to obtain sentence-level embeddings, which are processed by non-autoregressive transformer to predict pose latents in the disentangled latent space. A temporal mean-pooling supports length prediction, while the decoder is conditioned on time queries derived from an idle pose. Training uses a region-specific weighted L1 loss and KL term for regularization. The disentangled AE decoder then reconstructs regions keypoints to produce the final sign pose sequence.
@@ 3.1 Pose Autoencoder
To model the multimodal structure of sign language articulation, we design a structurally disentangled pose autoencoder that encodes and reconstructs sign poses within a compact latent space. Rather than aiming to achieve disentanglement in the strict unsupervised sense, we explicitly incorporated an architectural design that encourages the model to learn semantically organized representations. This design introduces a multi-modal latent factorization that aligns well with the physical structure and functional roles of each body part in sign language expression.
Our model processes 3D skeleton sequences comprising 8 upper body and 21 left hand, 21 right hand joints, along with 128 facial keypoints, each represented by (x, y, z) coordinates. As shown in Figure 1, the encoder is anatomically partitioned into four regions, upper body, right hand, left hand, and face, each handled by a dedicated linear encoder. The pose autoencoder takes these 3D joint coordinates arranged in a fixed, anatomically consistent order defined by the dataset’s skeleton topology (e.g., Mediapipe format).
Based on ablation results (Section 4.2), we fix the total latent space to 80 dimensions, assigning 16 to the face and distributing the remaining 64 across other regions in proportion to their joint counts; Upper body (8 joints): mapped from 8 × 3 to 8 dimensions, Right/Left hand (21 joints each): mapped from 21 × 3 to 28 dimensions per hand, Face (128 joints): mapped from 128 × 3 to 16 dimensions. Although the face has high raw input dimensionality, many of its keypoints—particularly in the cheeks, jawline, and forehead—exhibit correlated motion. This redundancy allows the encoder to use fewer latent dimensions for the face without significantly affecting reconstruction quality.
To accommodate the characteristics of each dataset, we use slightly different autoencoder architectures. For PHOENIX14T [3], each articulator is encoded using a single linear projection layer, allowing direct mapping from raw 3D joint inputs to their respective latent representations. This simple and efficient structure is sufficient for PHOENIX14T, which has a more constrained linguistic domain and relatively lower articulatory variability. The decoder modules also consist of single linear layers per region, mirroring the encoder structure. For CSL-Daily [36], which involves broader linguistic context and higher articulatory variability, we introduce an additional projection layer with nonlinearity for Right/Left hand and face regions. Specifically, each region is encoded using a two-layer MLP comprising a linear projection, a PReLU activation, and a second linear layer. This increased representational capacity helps stabilize the latent space, particularly for hand and face regions where variability is highest. The intermediate hidden dimension is set to 40 for both hands and 96 for the face. Decoder modules follow the same structure, using symmetric two-layer MLPs with matching intermediate widths.
@@ 3.2 Transformer Model
Following the latent pose representation learned by the pose autoencoder, the second part of our framework generates sign pose embeddings conditioned on spoken language text. We utilize a Transformer-based architecture that maps sentence-level text to the latent space learned by the pose autoencoder.
@@
Our model, as illustrated in Figure 2, uses a Transformer encoder-decoder architecture. Input text to the encoder is represented by 768-dimensional BERT-based word embeddings, which are linearly projected to 512-dimension to match the model’s internal dimensions. The encoder consists of 3 layers with 4 attention heads and 1024-dimensional feed-forward networks, along with positional encoding to retain temporal ordering of the text sequence.
The decoder follows a non-autoregressive structure to reduce error accumulation and predicts 80-dimensional latent vectors corresponding to the autoencoder’s space. It includes 6 layers with 8 attention heads and the same 1024-dimensional feed-forward size. Temporal dynamics are initialized using learned time queries derived from a fixed stationary reference pose, where both hands rest downward as when signers are not performing a sign. This pose is projected into the decoder space and expanded across all time steps to provide a consistent initialization for sequence generation.
@@
Ground truth 80-dimensional pose encodings are extracted using the pretrained autoencoder and serve as structured supervision for transformer training. These encodings provide structurally disentangled representations of manual (hands) and non-manual (face, upper body) articulators, guiding the model in learning mappings from text to pose. To ensure alignment with the autoencoder’s latent space, we incorporate L1 and KL-based regularization during training, encouraging accurate and coherent pose generation.
@@
We also incorporate a length predictor to dynamically adjust the length of the generated sign sequences based on the given input sentences. First, a temporal mean pooling is applied, averaging across the temporal dimension while ignoring padding positions. This sentence-level embedding from the encoder is passed through a linear layer with sigmoid activation to predict a normalized length ratio (0–1), which corresponds to the proportion of the maximum possible sequence length required to represent the given input sentence. The module is trained using supervision from ground-truth sequence lengths, promoting temporal alignment consistent with semantic load, in line with the prior work [35].
@@
The predicted latent sequence is decoded into 3D pose keypoints using the pretrained pose autoencoder’s decoder. This design enables efficient, end-to-end sign pose generation without the need for additional refinement or post-processing.
3.3 Training Procedure
Our training follows a two-stage strategy for mapping spoken language to continuous sign pose sequences. In the first stage, we train a structurally disentangled autoencoder to learn compact latent representations that separately encode manual (hands) and non-manual (face, upper body) features. In the second stage, a non-autoregressive Transformer is trained to predict these latent representations from word-level text embeddings of the input sentence, supervised by the ground-truth encodings from the pretrained autoencoder.
3.3.1 Training the Pose Autoencoder
The structurally disentangled autoencoder is trained to minimize a weighted reconstruction loss across four anatomical regions, body (B), right hand (RH), left hand (LH), and face (F) combined with L1 regularization applied only to encoder weights to promote sparsity:
$\mathcal{L}_{\text{ae}}=\sum_{R\in\{\text{B},\,\text{RH},\,\text{LH},\,\text{F}\}}w_{R}\frac{1}{N_{R}}\sum_{i=1}^{N_{R}}\|\hat{\mathbf{x}}_{R}^{(i)}-\mathbf{x}_{R}^{(i)}\|_{1}+\lambda\sum_{j\in\{\text{enc}\}}\|\mathbf{W}_{j}\|_{1}$ (1)
Here, $\mathbf{x}_{R}^{(i)}$ and $\hat{\mathbf{x}}_{R}^{(i)}$ denote the ground truth and reconstructed keypoints for region $R$ at frame $i$, and $N_{R}$ is the number of frames. The first term is a weighted L1 reconstruction loss, where the weights $w_{R}$ allow articulator-specific emphasis during training. The second term is an L1 regularization loss applied solely to the encoder weights, a strategy used in sparse autoencoders [27]. This preserves the decoder’s full capacity to reconstruct fine-grained spatial dynamics, which is particularly critical for manual articulators like fingers.
3.3.2 Training Generator Model
The training objective of our transformer model is divided into two phase to effectively balance the learning of accurate pose encodings and structured latent space distribution. In the first-phase, the model is trained to predict ground truth (GT) pose encodings obtained from the pretrained pose autoencoder. In this phase, we minimize an L1 loss between the predicted encodings for the transformer model and the obtained GT encodings.
$\mathcal{L}_{\text{phase1}}=\sum_{R\in\{\text{B},\text{RH},\text{LH},\text{F}\}}w_{R}\cdot\frac{1}{T}\sum_{t=1}^{T}\left\|\hat{z}_{t}^{(R)}-z_{t}^{(R)}\right\|_{1}+\left\|\hat{r}-r\right\|_{1}$ (2)
Here, $\hat{z}_{t}^{(R)}$ and $z_{t}^{(R)}$ denote the predicted and ground-truth latent embeddings for region $R$ at timestep $t$, and $w_{R}$ are region-specific weights. The second term supervises the predicted sequence length ratio $\hat{r}$ using the ground-truth $r$. We will refer to this model as DARSLP.
In the second phase of the training, we incorporate a KL divergence term to promote generalization and preserve the distributions of the structurally disentangled latent space. We refer to the resulting model as DARSLP-KL, which extends the training of the first-phase model by adding regularization on the channel distributions (Eqn. 3). In this context, during training, we enforce the predicted latent channel distributions to align with precomputed priors derived from ground truth encodings for each articulator region.
$\mathcal{L}_{\text{phase2}}=\mathcal{L}_{\text{phase1}}+\sum_{R\in\{\text{B},\text{RH},\text{LH},\text{F}\}}\sum_{c\in R}D_{\text{KL}}\left(\mathcal{N}(\mu^{(c)},\sigma^{(c)})\,\|\,\mathcal{N}(\mu_{\text{prior}}^{(c)},\sigma_{\text{prior}}^{(c)})\right)$ (3)
Here, $\mu^{(c)}$ and $\sigma^{(c)}$ denote the batch-wise mean and variance of predicted latent channels, while $\mu_{\text{prior}}^{(c)}$ and $\sigma_{\text{prior}}^{(c)}$ are prior statistics computed from the autoencoder’s training embeddings.